After initial setup on Thursday the game was running OK however by Sunday two faults had developed. Firstly, some of the sprites were corrupted that was likely a game PCB issue and secondly the monitor colour balance had become poor.
Given that some sprites were OK and some were bad, the address multiplexors seemed as good a place as any to start. I setup the comparator for 74LS157 and began working through them. IC's 1C, 1D and 1E had no issue but IC's 3A, 3B and 3C failed. Further, the comparator on IC 3A (LS157) caused the sprites to be correct. The comparator showed pin 1 was incorrect. Pin 1 is an input on the 74LS157 from signal "MPX", so the comparator shouldn't have been able to flag anything. Looking at IC 3A pin 1 on the scope showed the signal was bad (thus, the theory here is that the bad logic levels on the signal caused the comparator to misbehave). The signal is driven from IC 3H (LS32). Replacing IC 3A (LS32) fixed the game.
To better allow for "field repairs" at the show I'd decided to put together a "Show Spares Box" of spare game PCB's to swap in when boards fail. As part of that effort I selected a best-of-the-bunch (#006) set to work on. The initial state was that a garbled score was displayed followed a watchdog reset.
Testing with the Arduino ICT flagged CPU board IC 2A (2114) and IC 2C (2114) as bad. Replacing both fixed the boot crash and the game was running OK but without active audio (only speaker hiss).
Inspecting the audio board found pin 38 of IC 4L (6802) bent under the CPU rather than in the socket. Fixing the bent pin didn't change anything - still no audio. The speech LED was lighting and the 5200 data bus was running. Inspecting IC's 4H & 4G (8910) pins 4, 3 & 38 showed active audio output, suggesting that the no audio problem was in the pre-amplifier stage. It was whilst I was looking for the audio problem the board lost all video output :(
A cursory poke around with a scope found that all the video clocks were idle. There was a brief display on power on or off, but otherwise nothing. The horizontal video counters IC 3L & IC 4L (LS161) seemed to have good outputs. The vertical video counter IC 3N (LS161) outputs were idle. Replacing IC 3N brought the video back.
Moving back to the audio problem, I used a clip attached to a headphone to verify that there was audio at R114 and R3, the input to the power amplifier IC 2B (TDA1510). The amplifier was getting very hot despite the large heatsink attached to it. I suspected it bad, or at least worth swapping out to verify it. Replacing the power amplifier brought back all the audio with no further audio issues.
In the intervening period waiting for power amplifiers to arrive the game developed lines in the sprites. Using a video probe suggested IC's 5E & 6E (LS157) were working OK. The video probe on IC's 7D & 7E (2148) also seemed OK but on IC 7K & 7L (2148) the activity was much quieter. Inspecting the control lines ~CST1 & ~CST2 found ~CST2 active with or without sprites on screen whereas ~CST1 was idle when no sprites were displayed. I suspected a problem further upstream of ~CST1. Since the video generation on this game uses two almost identical circuits that handle alternate lines it was possible to compare the two sides. I discovered that when the screen is clear of sprites A6, pin 1 of IC 7D (2148), was active but A6, pin 1 of IC 7K (2148), was idle. Since the address bus for these RAM were generated from block of four 74LS161 (IC's 7G, 8G, 8H, 7H) I setup the comparator to check them. When sprites were active on screen the comparator flagged all four as bad, that seemed unlikely and more likely the comparator was incompatible with the circuit. However, when sprites were not active on screen the comparator flagged only IC 7H as bad. Replacing IC 7H (LS161) fixed the sprites.
After several hours soak testing on the bench the game ended up stuck in DIP switch test mode. I'd seen this issue before when either the 8255 or switch matrix circuit was faulty. Using the on screen display to verify the switches found switch bank 4I not working. The scope showed that pin 12 of IC 4H (8255) wasn't active. Replacing the 8255 restored pin 12 to active and the high nibble of 4I was now working properly but the low nibble was stuck low on all switches. I suspected a bad IC 3H (40097) and replacing IC 3H fixed all the dip switches. After several more hours of soak testing no further issues occurred :)
I'd had colour drive stability issues on this chassis before. Poking around with a scope found green instability on TR23 & TR20. TR35 was also unstable, though only on one side. Replacing TR35 (BC307) made no difference. I began to suspect bad capacitors or resistors (the latter already replaced in the prior cap kit and a resistor fault is very rare). Replacing TR20 (BC237) fixed the green instability.
The next problem was poor contrast. The grid control had no effect on the picture. The brightness was poor in general and the colour drive looked binary - full on or full off. Using a 10x probe on the scope to look at the gun drives, KB, KR & KG, found bad KG gun drive. On the neck PCB the green signal was off at TR29 (BC237). Replacing TR29 fixed the KG gun drive and the chassis appeared to be working properly.
Back in the cabinet the picture was re-adjusted, looked OK and ran OK for several hours :)